# Littlewood, Paley and Almost-Orthogonality: a theory well ahead of its time

### v1

@@ Abstract.
Littlewood–Paley theory began with the classic paper of Littlewood and Paley (J. E. Littlewood, R. E. A. C. Paley, Theorems on Fourier Series and Power Series. J. Lond. Math. Soc. (1), 6 (1931), 230–33). We discuss this paper and its impact from a historical perspective. We include an outline of the results in the paper and their subsequent significance in relation to developments over the last century, and set them into the context of the current state of the art in harmonic analysis and beyond.
@@ 1. Introduction
Euclidean geometry, and more generally the study of euclidean spaces, is founded on the notion of orthogonality. The same holds for the study of Hilbert spaces, the infinite-dimensional generalisations of euclidean spaces, of which $L^{2}$ is the prototypical example. The fundamental and beautiful spectral theorem for bounded linear self-adjoint operators on a Hilbert space is entirely founded upon the concept of orthogonality. However, many problems occurring in analysis, mathematical physics, geometry, number theory and PDE, inter alia, require consideration of $L^{p}$ spaces where $p\in(1,\infty)$ is an index different from $2$. Is there a practical and useful substitute for orthogonality in such a setting, which might for example lead to results with the same flavour as the spectral theorem does for the case $p=2$?
The foundational paper of Littlewood and Paley, [LP1], addresses this question. An entire mathematical culture and body of thought, going by the name Littlewood–Paley theory, can be traced back to to it. In the present article we review this paper, and sketch some of its ramifications over the last nearly 100 years. In a nutshell, the paper introduced the notion that ideas and techniques based on orthogonality can be developed and deployed in settings where one has no right to believe that they should even exist.
@@
As we have observed, the space $L^{2}$, and operators on $L^{2}$ spaces, are susceptible to the techniques of Hilbert spaces, and in particular, orthogonality is baked-in. Two functions $f$ and $g$ are orthogonal if and only if
@@
The paradigm emerging from this perspective is that one may take a function $f$ or an operator $T$, (which we shall assume for the moment is a convolution operator), and decompose it into orthogonal pieces $Q_{j}f$, or $Q_{j}T=TQ_{j}$, work with each piece separately, and then use Pythagoras’ theorem to re-assemble the pieces.
When confronted with a problem whose correct expression is not in the language of $L^{2}$-spaces, but instead of that of $L^{p}$-spaces, where $1<p<\infty$ and $p\neq 2$, we seem doomed to failure if we entertain the idea of trying to use orthogonality techniques when analysing it. The genius of Littlewood and Paley, as manifested in this paper [LP1], is encapsulated by the notion that for $f\in L^{p}$, by considering the quadratic expression
@@
appearing in (1) and its quantitative relationship to $f$, we may indeed bring ideas of orthogonality to bear on problems where the appearance of $L^{p}$ for $p\neq 2$ is essential.
The aim of this article is to give a taste of the story of these and related quadratic expressions, or square functions, the $L^{p}$ estimates they satisfy, and a hint of their application and influence in the wider context of almost-orthogonality, over the last nearly one hundred years since Littlewood–Paley theory first emerged in the paper [LP1].
Disclaimer and acknowledgements. What follows is a personal, subjective and in places rather impressionistic account, which reflects the tastes, interests and experience of its author. There is no claim (explicit or implicit) to comprehensiveness in either the text or the references. In particular there will be many authors whose work relies upon and is intimately bound up with Littlewood–Paley theory, and indeed advances it in significant ways, but who are not mentioned here; I can only apologise in advance for such omissions. We make no attempt to address the endpoint cases $p=1$ and $p=\infty$, still less the cases $0<p<1$. (These too have a very rich history, bound up with Hardy spaces $H^{p}$ in their classical and real-variable incarnations, and the John–Nirenberg space BMO of functions of bounded mean oscillation which is the dual space of the real Hardy space $H^{1}$. See [SteinBig].) For a complementary perspective focusing on Zygmund’s work on square functions see [SteinZygmund]; that article gives an authoritative and thourough account of developments up to the early 1980s. See also [CarberyCZsurvey] for an account of related developments up to 1993. The author would like to thank Jonathan Hickman for comments on an earlier version of the paper, and Joseph Feneuil, Linhan Li and Michele Villa for assistance with Section 8. The comments of the referee have led to many presentational improvements. All errors, omissions and misrepresentations remaining are the sole responsibility of the author.
@@ 2. What’s in the Littlewood–Paley paper (and what’s not)
In this paper [LP1], the programme of bringing orthogonality techniques to bear on $L^{p}$ for $p\neq 2$ is addressed in the specific setting of $L^{p}(\mathbb{T})$ for $1<p<\infty$. Here, $\mathbb{T}=\mathbb{R}/\mathbb{Z}$ is the unit torus which we can identify with (1/2,1/2](-1/2,1/2], and the Fourier transform $f\mapsto\widehat{f}$ is the Fourier coefficient mapping from $L^{2}(\mathbb{T})$ to $\ell^{2}(\mathbb{Z})$ given by
@@
For each $1<p<\infty$, there are constants $A_{p}$ and $B_{p}$ such that
@@
Note that this result immediately implies that for a lacunary trigonometric series $\sum_{j\geq 0}a_{j}e^{2\pi i2^{j}x}$, its $L^{p}$ norms for $1<p<\infty$ are all comparable to $(\sum_{j\geq 0}|a_{j}|^{2})^{1/2}$, since each dyadic block $E_{j}$ then contains at most one non-zero Fourier frequency. The “digitalised" version of this consequence – in which the term $\sin(2^{j+1}\pi x)$ implicitly appearing in a lacunary trigonometric series is replaced by its signum – goes back a little further to Khintchine [MR1544623]. Results such as this capture the concept of “square-root cancellation" which serves as an intuitional lynchpin across modern harmonic analysis and beyond.
As a pretty much immediate consequence of Theorem 2.1, Littlewood and Paley state:
@@
For each $1<p<\infty$, there is a constant $C_{p}$ such that if the numerical sequence $\{r_{j}\}$ satisfies $|r_{j}|\leq 1$ for all $j$, then
@@
Interestingly, in [LP1] this result is stated under the stronger hypothesis that $|r_{j}|=1$ for all $j$, and, in that case, the stronger conclusion that $\|\sum_{j\geq 0}r_{j}Q_{j}f\|_{p}\approx\|f\|_{p}$ is not mentioned. Here as below, the symbol $\approx$ indicates that the ratio of the left and right hand sides is bounded above and below by constants depending only on $p$ (and the underlying dimension when that enters the story later).
The proof of Theorem 2.2 relies upon applications of both inequalities in (2) and can be summarised as the chain of inequalities
$\|\sum_{j\geq 0}r_{j}Q_{j}f\|_{p}\leq A_{p}^{-1}\|\big(\sum_{k\geq 0}|Q_{k}(\sum_{j\geq 0}r_{j}Q_{j}f)|^{2}\big)^{1/2}\|_{p}\leq A_{p}^{-1}\|\big(\sum_{j\geq 0}|Q_{j}f|^{2}\big)^{1/2}\|_{p}\leq A_{p}^{-1}B_{p}\|f\|_{p},$
since $Q_{k}Q_{j}=0$ for $j\neq k$ and $Q_{j}^{2}=Q_{j}$.
Theorem 2.2 is a primordial version of later Fourier multiplier theorems going by the names of Marcinkiewicz [^2] In this case, its ultimate formulation is still a matter of current study, see for example [bakasetal]. and Hörmander–Mikhlin, extensive discussion of both of which can be found in [Stein]. In fact, the proofs of these more general multiplier theorems are not much more difficult than that of Theorem 2.2.
@@
where, for each $1<r<\infty$, $w\mapsto\mathcal{M}w$ is some (possibly non-linear) non-negative operator which is bounded on $L^{r}(\mathbb{T})$, say with constant $K_{r}$. (We will discuss verification of this hypothesis below.) We shall then be able to conclude that
@@
which establishes the desired conclusion under the hypothesis (3). A very similar argument also shows that
$\|\big(\sum_{j\geq 0}|T_{j}f|^{2}\big)^{1/2}\|_{p}\leq C_{p}^{\prime}\|f\|_{p}$
under the same hypothesis (3).
It is perhaps more traditional to see the above argument presented as a pointwise inequality of the form
@@
where $T=\sum_{j\geq 0}T_{j}$, and $\mathcal{Q}_{1}$ and $\mathcal{Q}_{2}$ are certain variants of $\mathcal{Q}$, enjoying the same $L^{p}$ inequalities as $\mathcal{Q}$ when $p\geq 2$. (Here and below we use the notation $\lesssim$ to denote the existence of a constant bounding the ratio of the left hand side to the right hand side.)
This argument crystallises the idea that the Littlewood–Paley theory allows us to derive conclusions about a sum of operators $\sum_{j}T_{j}$ from detailed uniform knowledge of the $T_{j}$ individually, as in (3).
Condition (3) is easily verified under the hypotheses [^3] These are essentially quantitative uniform smoothness hypotheses on the Fourier multipliers $m_{j}$. of the Hörmander–Mikhlin multiplier theorem, for example. In that case we simply take $\mathcal{M}w$ to be (a constant multiple of) the Hardy–Littlewood maximal function of $w$. (Incidentally, the hypothesis (3) can be significantly weakened, see [CarbCZ] and [Seeger].) We turn next to the circle of ideas around maximal functions.
In [LP1], Littlewood and Paley were strongly motivated by potential applications to Lusin’s problem concerning pointwise convergence of Fourier series. While this was settled some thirty-five years later by Carleson [Carleson] and Hunt [MR238019], in the early 1930’s the question was still very much alive. Already it was understood that the correct way to control pointwise convergence behaviour in a quantitative way is with the use of appropriate maximal functions, of which the newly-born Hardy–Littlewood maximal function [MR1555303] was and is the example par excellence. The (one-dimensional) Hardy–Littlewood maximal operator
@@
One seeks to prove that (either version of) $S_{\ast}f$ is finite almost everywhere for all $f$ in as large a subclass $\mathcal{L}$ of $L^{1}$ as possible – then pointwise almost-everywhere convergence for the same class will hold, via the fact that it always holds for sufficiently nice functions $f$ by classical results, and then a density argument. Kolmogorov [zbMATH02586338] had shown that there exist functions in $L^{1}$ whose Fourier series diverge almost everywhere, and indeed everywhere, but the question for $\mathcal{L}=L^{p}$ for $p>1$ was still open. And the best way to obtain almost-everywhere finiteness of $S_{\ast}f$ for $f\in\mathcal{L}\subseteq L^{p}$ is to show that we have an a priori inequality of the form
@@
In this paper [LP1] Littlewood and Paley announce:
@@
We need to say a few words about the condition on $f$ in part (ii). It is to be thought of as a rather mild smoothness condition on $f$. This is because if $f$ and $f^{\prime}$ are continuous, then $\widehat{f^{\prime}}(k)=-2\pi ik\widehat{f}(k)$, as an elementary calculation shows. Similarly, for $m\geq 1$, $\widehat{f^{(m)}}(k)=(-2\pi ik)^{m}\widehat{f}(k)$. Moreover the $L^{p}$ norms of the functions with Fourier coefficients $(-2\pi ik)^{m}\widehat{f}(k)\mbox{ and }(2\pi|k|)^{m}\widehat{f}(k)$ are equivalent, because they are related by a Fourier multiplier of the type described directly after Theorem 2.2. Defining the fractional derivative $D^{s}f$ of $f$ (where $s$ is now any non-negative real number) as the function (if it exists!) whose Fourier coefficients are $(2\pi|k|)^{s}\widehat{f}(k)$, it is thus natural to think of $\widehat{f}(j)\log(|j|+2)^{1/p}$ as representing the Fourier coefficients of a logarithmic variant (near $s=0$) of the fractional derivative $D^{s}f$. [^4] Indeed, the Sobolev space $L^{p}_{s}$ can be defined to consist of those $f\in L^{p}$ such that $D^{s}f\in L^{p}$.
Part (i) of this theorem in the case $p=2$ had already been observed by Hardy and Littlewood. It also holds for $2<p<\infty$, but the pointwise convergence consequence in that case already follows from the case $p=2$. The pointwise convergence consequence of part (ii) in the case $p=2$ had been observed by Kolmogorov, Seliverstov and Plessner, but the maximal function bound was new both in this case and the case $1<p<2$. There is a weaker bound given in the paper for the maximal function in part (ii) when $p>2$, but Littlewood and Paley were clearly dissatisfied with it even at the time. (Nevertheless it was not until Carleson’s work in the 1960’s that even that weaker assertion received any quantitative strenghtening whatsoever.)
The proof of part (i) of this theorem is rather straightforward, but has been very influential. The idea, which goes back to Kaczmarz and Zygmund, [^5] Stein [SteinZygmund] attributes the first appearance of quadratic expressions in harmonic analysis, albeit in the auxiliary role we are now discussing, to Kaczmarz and Zygmund in 1926, though this was perhaps not realised at the time by Hardy, Littlewood and Paley. is as follows: we can decompose $S_{2^{k}}$ into contributions $P_{k}$ (an averaging operator at scale $2^{-k}$) for frequencies of size at most $2^{k}$, and $T_{k}$ for frequencies of size about $2^{k}$ in such a way that
$S_{2^{k}}=P_{k}+T_{k}$
where we have an easy pointwise control $\sup_{k}|P_{k}f(x)|\lesssim Mf(x)$, and where $\{T_{k}\}$ satisfies the considerations of the discussion after Theorem 2.2. The key point is then that
@@
This argument serves as the prototype of a whole slew of applications of Littlewood–Paley theory to the analysis of maximal functions. The thrust of the method is that a quadratic expression mediates between – or acts as a Tauberian condition linking – the more difficult maximal function under study, and a more basic maximal function which is already understood. This reduces matters to establishing $L^{p}$ boundedness of the mediating quadratic expression. Hitherto, the study of maximal averaging operators in higher dimensions had relied upon establishing delicate geometric covering lemmas somewhat analogous to the Vitali or Wiener covering lemmas for balls. But in many cases of interest, such as spherical averages (in which averages over balls appearing in the higher dimensional Hardy–Littlewood maximal operator are replaced by averages over spheres, see [SteinBig]), it was not known how to do this. Mediation by quadratic expressions offered an entirely different perspective on such problems, even in the case $p=2$, and in the 1970’s and 1980’s led to spectacular success not only for the spherical maximal operator but also for (amongst many others) the study of differentiation in lacunary directions [NSW]. The more basic, “already-understood" maximal functions in these two cases are the classical Hardy–Littlewood maximal operator and the Hardy–Littlewood maximal operator with parabolic dilations (see below) respectively.
While Theorem 2.3 has been superseded in strict terms by the foundational work of Carleson and Hunt [Carleson], [MR238019], the fundamental ideas behind it continue to remain highly influential to this day.
The paper under discussion contains no proofs, and no details of the application of Theorem 2.1 to Theorem 2.3 part (ii). These are given in the subsequent papers [LP2] and [LP3] which appeared after Paley’s tragic death in a skiing accident in 1933. The paper does sketch, in broad strokes, the main lines of argument, which are inextricably rooted in the function theory of one complex variable, and which rely on various considerations which are available only in that setting. It makes implicit contact with other quadratic expressions such as the area function introduced by Lusin in 1930, and later explored by Marcinkiewicz and Zygmund.

### v2

@@ Abstract.
The classic paper J. E. Littlewood, R. E. A. C. Paley, Theorems on Fourier Series and Power Series. J. Lond. Math. Soc. (1), 6 (1931), 230–33, marked the birth of Littlewood–Paley theory. We discuss this paper and its impact from a historical perspective. We include an outline of the results in the paper and their subsequent significance in relation to developments over the last century, and set them into the context of the current state of the art in harmonic analysis and beyond.
@@ 1. Introduction
Euclidean geometry, and more generally the study of euclidean spaces, is rooted in the notion of orthogonality. The same holds for the study of Hilbert spaces, the infinite-dimensional generalisations of euclidean spaces, of which $L^{2}$ is the prototypical example. The fundamental and beautiful spectral theorem for bounded linear self-adjoint operators on a Hilbert space is entirely founded upon the concept of orthogonality. However, many problems occurring in analysis, mathematical physics, geometry, number theory and PDE, inter alia, require consideration of $L^{p}$ spaces, where $p\in(1,\infty)$ is an index different from $2$. Is there a practical and useful substitute for orthogonality in such a setting, which might for example lead to results with flavours similar to those that the spectral theorem yields in the case $p=2$?
The foundational paper of Littlewood and Paley, [50], begins to address this question. An entire mathematical culture and body of thought, now going by the name Littlewood–Paley theory, can be traced back to to it. In the present article we review this paper, and sketch some of its ramifications over the last nearly 100 years. In a nutshell, the paper introduced the notion that ideas and techniques based on orthogonality can be developed and deployed in settings where one has no right to believe that they should even exist.
@@
As we have observed, orthogonality is baked into the Hilbert space $L^{2}$, and to operators on $L^{2}$ spaces. Two functions $f$ and $g$ in $L^{2}$ are orthogonal if and only if
@@
The philosophy which emerges from this perspective is that one may take a function $f$ or an operator $T$, (which we shall assume for the moment is a convolution operator), and decompose it into orthogonal pieces $Q_{j}f$, or $Q_{j}T=TQ_{j}$, work with each piece separately, and then use Pythagoras’ theorem to re-assemble the pieces.
When confronted instead with a problem whose correct expression is not in the language of $L^{2}$-spaces, but instead of that of $L^{p}$ spaces, where $1<p<\infty$ and $p\neq 2$, we seem doomed to failure if we entertain the idea of trying to analyse it using techniques of orthogonality. The genius of Littlewood and Paley, as manifested in this paper [50], is encapsulated in the notion that for $f\in L^{p}$, by considering the quadratic expression
@@
which appears in (1), and its quantitative relationship to $f$, we may indeed bring ideas of orthogonality to bear on problems which are essentially $L^{p}$ in nature.
The aim of this article is to give a taste of the story of these and related quadratic expressions, or square functions, the $L^{p}$ estimates they satisfy, and a hint of their application and influence in the wider context of almost-orthogonality, over the last nearly one hundred years since Littlewood–Paley theory first emerged in [50].
Disclaimer and acknowledgements. What follows is a personal, subjective and in places rather impressionistic account, which reflects the tastes, interests and experience of its author. There is no claim (explicit or implicit) to comprehensiveness in either the text or the references. In particular there will be many authors whose work relies upon and is intimately bound up with Littlewood–Paley theory, and which indeed advances it in significant ways, but who are not mentioned here; I can only apologise in advance for such omissions. We make no attempt to address the endpoint cases $p=1$ and $p=\infty$, still less the cases $0<p<1$. (These too have a very rich history, bound up with Hardy spaces $H^{p}$ in their classical and real-variable incarnations, and the John–Nirenberg space BMO of functions of bounded mean oscillation which is the dual space of the real Hardy space $H^{1}$. See [71].) For a complementary perspective focusing on Zygmund’s work on square functions see [67]; that article gives an authoritative and thorough account of matters up to the early 1980s. See also [17] for an account of related developments up to 1993. The author would like to thank Jonathan Hickman for comments on an earlier version of the paper, and Joseph Feneuil, Linhan Li and Michele Villa for guidance on Section 8. The comments of the referee have led to many presentational improvements. All errors, omissions and misrepresentations remaining are the sole responsibility of the author.
@@ 2. What’s in the Littlewood–Paley paper (and what’s not)
In this paper [50], the programme of bringing orthogonality techniques to bear on $L^{p}$ for $p\neq 2$ is addressed in the specific setting of $L^{p}(\mathbb{T})$ for $1<p<\infty$. Here, $\mathbb{T}=\mathbb{R}/\mathbb{Z}$ is the unit torus or circle group which we can identify with (1/2,1/2](-1/2,1/2], and the Fourier transform $f\mapsto\widehat{f}$ is the Fourier coefficient mapping from $L^{2}(\mathbb{T})$ to $\ell^{2}(\mathbb{Z})$ given by
@@
For each $1<p<\infty$, there are constants $A_{p}$ and $B_{p}$, such that for all $f\in L^{p}$,
@@
Note that this result immediately implies that for a lacunary trigonometric series $\sum_{j\geq 0}a_{j}e^{2\pi i2^{j}x}$, its $L^{p}$ norms for $1<p<\infty$ are all comparable to $(\sum_{j\geq 0}|a_{j}|^{2})^{1/2}$, since each dyadic block $E_{j}$ then contains at most one non-zero Fourier frequency. The “digitalised" version of this consequence – in which the term $\sin(2^{j+1}\pi x)$ implicitly appearing in a lacunary trigonometric series is replaced by its signum – goes back a little further to Khintchine [47]. Results such as this capture the concept of “square-root cancellation" which serves as an intuitional lynchpin across modern harmonic analysis and beyond.
As a pretty much immediate consequence of Theorem 2.1, Littlewood and Paley derive:
@@
For each $1<p<\infty$, there is a constant $C_{p}$ such that if the numerical sequence $\{r_{j}\}$ satisfies $|r_{j}|\leq 1$ for all $j$, then, for all $f\in L^{p}$,
@@
Interestingly, in [50] this result is stated under the stronger hypothesis that $|r_{j}|=1$ for all $j$, and, in that case, the stronger conclusion that $\|\sum_{j\geq 0}r_{j}Q_{j}f\|_{p}\approx\|f\|_{p}$ is not mentioned. Here as below, the symbol $\approx$ indicates that the ratio of the left and right hand sides is bounded above and below by constants depending only on $p$ (and the underlying dimension when that enters the story later).
The proof of Theorem 2.2 relies upon applications of both inequalities in (2), and can be summarised by the chain of inequalities
$\|\sum_{j\geq 0}r_{j}Q_{j}f\|_{p}\leq A_{p}^{-1}\|\big(\sum_{k\geq 0}|Q_{k}(\sum_{j\geq 0}r_{j}Q_{j}f)|^{2}\big)^{1/2}\|_{p}\leq A_{p}^{-1}\|\big(\sum_{k\geq 0}|Q_{k}f|^{2}\big)^{1/2}\|_{p}\leq A_{p}^{-1}B_{p}\|f\|_{p},$
holding since the projection operators $Q_{k}$ satisfy $Q_{k}^{2}=Q_{k}$ and $Q_{k}Q_{j}=0$ for $j\neq k$.
Theorem 2.2 is a primordial version of subsequently established Fourier multiplier theorems going by the names of Marcinkiewicz [^2] In this case, its ultimate formulation is still a matter of current study, see for example [1]. and Hörmander–Mikhlin, extensive discussion of both of which (in the euclidean case) can be found in [69]. In fact, the proofs of these more general multiplier theorems are not much more difficult than that of Theorem 2.2, as we now demonstrate.
@@
where, for each $1<r<\infty$, $w\mapsto\mathcal{M}w$ is some (possibly non-linear) non-negative operator which is bounded on $L^{r}(\mathbb{T})$, say with constant $K_{r}$. (We will discuss verification of this hypothesis below. [^3] The hypothesis (3) can be significantly weakened, see [16] and [66].) We shall then be able to conclude that
@@
which establishes the desired conclusion under the hypothesis (3).
Remark.
(i) A very similar argument shows that under the same hypothesis (3) we also have $\|\big(\sum_{j\geq 0}|T_{j}f|^{2}\big)^{1/2}\|_{p}\leq C_{p}^{\prime}\|f\|_{p}$.
(ii) It is perhaps more traditional to see the above argument presented as a pointwise inequality of the form
@@
where $T=\sum_{j\geq 0}T_{j}$, and $\mathcal{Q}_{1}$ and $\mathcal{Q}_{2}$ are certain variants of $\mathcal{Q}$, enjoying the same $L^{p}$ inequalities as $\mathcal{Q}$ when $p\geq 2$, see [69]. (Here and below we use the notation $\lesssim$ to denote the existence of a constant bounding the ratio of the left hand side to the right hand side.)
(iii) This argument crystallises the idea that the Littlewood–Paley theory allows us to derive conclusions about a sum of operators $\sum_{j}T_{j}$ from detailed uniform knowledge of the $T_{j}$ individually, as in (3).
The uniform weighted assumption (3) is easily verified, for example, under the hypotheses [^4] These are essentially quantitative uniform smoothness hypotheses on the Fourier multipliers $m_{j}$. of the Hörmander–Mikhlin multiplier theorem. In that case we simply take $\mathcal{M}w$ to be (a constant multiple of) the Hardy–Littlewood maximal function of $w$. We turn next to the circle of ideas around maximal functions.
In [50], Littlewood and Paley were strongly motivated by potential applications to Lusin’s problem concerning pointwise convergence of Fourier series. While this was settled some thirty-five years later by Carleson [19] and Hunt [44], in the early 1930’s the question was still very much alive. Already it was understood that the correct way to control pointwise convergence behaviour in a quantitative way is with the use of appropriate maximal functions, of which the newly-born Hardy–Littlewood maximal function [41] was and is the example par excellence. The (one-dimensional) Hardy–Littlewood maximal operator
@@
One seeks to prove that (either version of) $S_{\ast}f$ is finite almost everywhere for all $f$ in as large a subclass $\mathcal{L}$ of $L^{1}$ as possible – then pointwise almost-everywhere convergence for the same class will hold, via the fact that it always holds for sufficiently nice functions $f$ by classical results, and then a density argument. Kolmogorov [48] had shown that we cannot take $\mathcal{L}$ to be all of $L^{1}$ since there exist functions in $L^{1}$ whose Fourier series diverge almost everywhere, and indeed everywhere, but the question for $\mathcal{L}=L^{p}$ for $p>1$ was still open. And the best way to obtain almost-everywhere finiteness of $S_{\ast}f$ for $f\in\mathcal{L}\subseteq L^{p}$ is to show that we have an a priori inequality of the form
@@
In this paper [50] Littlewood and Paley announce:
@@
We need to say a few words about the condition on $f$ in part (ii). It is to be thought of as a rather mild smoothness condition on $f$. This is because if $f$ and $f^{\prime}$ are continuous, then $\widehat{f^{\prime}}(k)=-2\pi ik\widehat{f}(k)$, as an elementary calculation shows. Similarly, for $m\geq 1$, $\widehat{f^{(m)}}(k)=(-2\pi ik)^{m}\widehat{f}(k)$. Moreover the $L^{p}$ norms of the functions with Fourier coefficients $(-2\pi ik)^{m}\widehat{f}(k)\mbox{ and }(2\pi|k|)^{m}\widehat{f}(k)$ are equivalent, because they are related by a Fourier multiplier of Hörmander–Mikhlin type as described after Theorem 2.2. Defining the fractional derivative $D^{s}f$ of $f$ (where $s$ is now any non-negative real number) as the function (if it exists!) whose Fourier coefficients are $(2\pi|k|)^{s}\widehat{f}(k)$, it is thus natural to think of $\widehat{f}(j)\log(|j|+2)^{1/p}$ as representing the Fourier coefficients of a logarithmic variant (near $s=0$) of the fractional derivative $D^{s}f$. [^5] Indeed, the Sobolev space $L^{p}_{s}$ can be defined to consist of those $f\in L^{p}$ such that $D^{s}f\in L^{p}$.
Part (i) of this theorem in the case $p=2$ had already been observed by Hardy and Littlewood. It also holds for $2<p<\infty$, but the pointwise convergence consequence in that case already follows from the case $p=2$. The pointwise convergence consequence of part (ii) in the case $p=2$ had been observed by Kolmogorov, Seliverstov and Plessner, but the maximal function bound was new both in this case and the case $1<p<2$. The bound for the maximal function in part (ii) when $p>2$ is also given in the paper, but under a stronger than desired hypothesis on the Fourier coefficients, and Littlewood and Paley were clearly dissatisfied with it even at the time. Nevertheless, this stronger hypothesis was weakened (and indeed removed altogether) only with Carleson’s work in the 1960’s.
The proof of part (i) of this theorem is rather straightforward, but it has been enormously influential. The idea, which goes back to Kaczmarz and Zygmund, [^6] Stein [67] attributes the first appearance of quadratic expressions in harmonic analysis, albeit in the auxiliary role we are now discussing, to Kaczmarz and Zygmund in 1926, though this was perhaps not known at the time by Hardy, Littlewood and Paley. is as follows: we can decompose $S_{2^{k}}$ into contributions $P_{k}$, which is an averaging operator at scale $2^{-k}$, corresponding to frequencies of size at most $2^{k}$, and $T_{k}$ for frequencies of size about $2^{k}$, in such a way that
$S_{2^{k}}=P_{k}+T_{k},$
where we have an easy pointwise control $\sup_{k}|P_{k}f(x)|\lesssim Mf(x)$, and where $\{T_{k}\}$ satisfies the considerations of Remark (i) after Theorem 2.2. The key point is then that
@@
This argument serves as the prototype of a vast array of applications of Littlewood–Paley theory to the analysis of maximal functions. The thrust of the method is that a quadratic expression mediates between – or acts as a Tauberian condition linking – the more difficult maximal function under study, and a more basic maximal function which is already understood. This reduces matters to establishing $L^{p}$ boundedness of the mediating quadratic expression. Hitherto, the study of maximal averaging operators in higher dimensions had relied upon establishing delicate geometric covering lemmas somewhat analogous to the well-known Vitali or Wiener covering lemmas for balls. But in many cases of interest, such as spherical averages (in which averages over balls appearing in the higher dimensional Hardy–Littlewood maximal operator are replaced by averages over spheres, see [71]), it was not known how to do this. Mediation by quadratic expressions offered an entirely different perspective on such problems, even in the case $p=2$, and in the 1970’s and 1980’s led to spectacular success not only for the spherical maximal operator but also for (amongst many others) the study of differentiation in lacunary directions [60]. The more basic, “already-understood" maximal functions in these two particular cases are the classical Hardy–Littlewood maximal operator and the Hardy–Littlewood maximal operator with parabolic dilations (see below) respectively.
While Theorem 2.3 has been superseded in strict terms by the seminal work of Carleson and Hunt [19], [44], the fundamental ideas behind it continue to remain highly influential to this day.
The paper under discussion contains no proofs, and no details of the application of Theorem 2.1 to Theorem 2.3 part (ii). These are given in the subsequent papers [51] and [52] which appeared after Paley’s tragic death in a skiing accident in 1933. The paper does sketch, in broad strokes, the main lines of argument, which are inextricably rooted in the function theory of one complex variable, and which rely on various considerations which are available only in that setting. It makes implicit contact with other quadratic expressions such as the area function introduced by Lusin in 1930, and later explored by Marcinkiewicz and Zygmund.